# DAVE: A VLM Vision Encoder for Document Understanding and Web Agents

### v2

@@ 5.4 Ablations
Image Inter-Patch Variance. To show the instability of the MAE training on document and web image discussed in Section 3.3, we plot the distribution density of inter-patch variance across different data sources in Figure 3. Compared to ImageNet (deng2009imagenet), document and web images exhibit much lower variance. This distributional gap underscores the need for a specialized vision encoder capable of handling such data. Additional implementation details are provided in Appendix E.
@@
Different Merge Setting. As described in Section 3.3, we use weight merging to create a decoder-agnostic vision encoder. Table 4(a) examines the effect of different merge strategies. Our end-to-end approach for learning merge coefficients outperforms both simple averaging and the heuristic-based Fisher Merge (matena2022merging) method, which weights parameters by their estimated importance using the Fisher information. Table 4(b) evaluates merging DAVE from different LLM backbones, where performance improves consistently with the number of merged encoders.
@@ B.1 Implementation Details
We use the official LLaVA-Next (li2024llava) repo with some customization to support different LLM and encoder settings, and we use 32 H200 GPUs for the training. To handle high resolution images, we use AnyRes image tilting with size of 384x384. The hyperparameters are listed in Table 7.
@@ B.2 Datasets
We use a wide range of publicly available data for our supervised pretraining. Notably, all the data excluding UGround are reformulated into layout and information extraction. For more detail on the data curation, refer to Granite Vision (team2025granite). For the Self-curated Data, we use PaddleOCR as our OCR model. We first filter arXiv papers to include only those published in venues or journals, ensuring a high-quality source corpus. During post-processing, we remove all bounding boxes with confidence scores below 0.5 to ensure accurate OCR localization. We then randomly resize, crop, and add noise to prevent the model from overfitting to a specific PDF size. Next, we normalize all bounding box coordinates to the range [0, 999] with respect to the new size. To control task difficulty, we discard bounding boxes that contain either too few words (e.g., single tokens) or excessively long paragraphs. Finally, from each remaining (category, bounding box, text) tuple, we construct diverse training examples for bbox-to-text, text-to-bbox, and bbox-to-category tasks. The data composition is shown in Table 8.
@@ B.3 Weight-Merging
For the weight-merging, we randomly sample 10k document and web images from self-supervised pretraining as our training data. We use the AdamW Optimizer (loshchilov2017decoupled), a learning rate of 1e-4, and a batch size of 32 to train for 20 epoches. The only trainable component is the newly introduced merge coefficient, while all the vision encoders remain frozen.
@@ C.1 DocBank
Implementation. DocBank consists of 13 categories. For each category, we finetune the vision encoder with an attention pooling network on the corresponding training set, and perform evaluation on the validation set. Before training, each image is processed and resized according to the default preprocessor for the specific vision encoder. For each category, we train with 5 epoch, a learning rate of 1e-5, weight decay of 1e-4, and a batch size of 32. We use Smooth L1 objective (huber1964robust) for the training. The per-category result is shown in Table 9.
@@ C.2 Doclaynet
Implementation. Doclaynet consists of 11 distinct classes, and in our implementation, we introduce an additional class for the area with no class label. We finetune the vision encoder along with an MLP projector to project the visual feature patches into a segmentation mask. Then we use crossentropy (shannon1948mathematical) between the prediction and the ground truth segmentation as objective and train on the training set. We train for 2 epoch, with a learning rate of 1e-5, weight decay of 1e-4, and a batch size of 32.

### v3

@@ 5.4 Ablations
Image Inter-Patch Variance. To show the instability of the MAE training on document and web image discussed in Section 3.3, we plot the distribution density of inter-patch variance across different data sources in Figure 3. Compared to ImageNet (Deng et al., 2009), document and web images exhibit much lower variance. This distributional gap underscores the need for a specialized vision encoder capable of handling such data. Additional implementation details are provided in Appendix E.
@@
Different Merge Setting. As described in Section 3.3, we use weight merging to create a decoder-agnostic vision encoder. Table 4(a) examines the effect of different merge strategies. Our end-to-end approach for learning merge coefficients outperforms both simple averaging and the heuristic-based Fisher Merge (Matena & Raffel, 2022) method, which weights parameters by their estimated importance using the Fisher information. Table 4(b) evaluates merging DAVE from different LLM backbones, where performance improves consistently with the number of merged encoders.
@@ B.1 Implementation Details
We use the official LLaVA-Next (Li et al., 2024) repo with some customization to support different LLM and encoder settings, and we use 32 H200 GPUs for the training. To handle high resolution images, we use AnyRes image tilting with size of 384x384. The hyperparameters are listed in Table 7.
@@ B.2 Datasets
We use a wide range of publicly available data for our supervised pretraining. Notably, all the data excluding UGround are reformulated into layout and information extraction. For more detail on the data curation, refer to Granite Vision (Team et al., 2025). For the Self-curated Data, we use PaddleOCR as our OCR model. We first filter arXiv papers to include only those published in venues or journals, ensuring a high-quality source corpus. During post-processing, we remove all bounding boxes with confidence scores below 0.5 to ensure accurate OCR localization. We then randomly resize, crop, and add noise to prevent the model from overfitting to a specific PDF size. Next, we normalize all bounding box coordinates to the range [0, 999] with respect to the new size. To control task difficulty, we discard bounding boxes that contain either too few words (e.g., single tokens) or excessively long paragraphs. Finally, from each remaining (category, bounding box, text) tuple, we construct diverse training examples for bbox-to-text, text-to-bbox, and bbox-to-category tasks. The data composition is shown in Table 8.
@@ B.3 Weight-Merging
For the weight-merging, we randomly sample 10k document and web images from self-supervised pretraining as our training data. We use the AdamW Optimizer (Loshchilov & Hutter, 2017), a learning rate of 1e-4, and a batch size of 32 to train for 20 epoches. The only trainable component is the newly introduced merge coefficient, while all the vision encoders remain frozen.
@@ C.1 DocBank
Implementation. DocBank consists of 13 categories. For each category, we finetune the vision encoder with an attention pooling network on the corresponding training set, and perform evaluation on the validation set. Before training, each image is processed and resized according to the default preprocessor for the specific vision encoder. For each category, we train with 5 epoch, a learning rate of 1e-5, weight decay of 1e-4, and a batch size of 32. We use Smooth L1 objective (Huber, 1964) for the training. The per-category result is shown in Table 9.
@@ C.2 Doclaynet
Implementation. Doclaynet consists of 11 distinct classes, and in our implementation, we introduce an additional class for the area with no class label. We finetune the vision encoder along with an MLP projector to project the visual feature patches into a segmentation mask. Then we use crossentropy (Shannon, 1948) between the prediction and the ground truth segmentation as objective and train on the training set. We train for 2 epoch, with a learning rate of 1e-5, weight decay of 1e-4, and a batch size of 32.